\begin{thebibliography}{99}

%\cite{Unruh:1976db}
\bibitem{Unruh:1976db} 
  W.~G.~Unruh,
  ``Notes on black hole evaporation,''
  Phys.\ Rev.\ D {\bf 14}, 870 (1976).
  doi:10.1103/PhysRevD.14.870
  %%CITATION = doi:10.1103/PhysRevD.14.870;%%
  %2784 citations counted in INSPIRE as of 12 Mar 2019

%\cite{Castorina:2012yg}
\bibitem{Castorina:2012yg} 
  P.~Castorina and M.~Finocchiaro,
  ``Symmetry Restoration By Acceleration,''
  J.\ Mod.\ Phys.\  {\bf 3}, 1703 (2012)
  doi:10.4236/jmp.2012.311209
  [arXiv:1207.3677 [hep-th]].
  %%CITATION = doi:10.4236/jmp.2012.311209;%%
  %9 citations counted in INSPIRE as of 03 Apr 2019
  
%\cite{Ohsaku:2004rv}
\bibitem{Ohsaku:2004rv} 
  T.~Ohsaku,
  ``Dynamical chiral symmetry breaking and its restoration for an accelerated observer,''
  Phys.\ Lett.\ B {\bf 599}, 102 (2004)
  doi:10.1016/j.physletb.2004.08.019
  [hep-th/0407067].
  %%CITATION = doi:10.1016/j.physletb.2004.08.019;%%
  %20 citations counted in INSPIRE as of 03 Apr 2019
  
%\cite{Takeuchi:2015nga}
\bibitem{Takeuchi:2015nga} 
  S.~Takeuchi,
  ``Bose–Einstein condensation in the Rindler space,''
  Phys.\ Lett.\ B {\bf 750}, 209 (2015)
  doi:10.1016/j.physletb.2015.09.013
  [arXiv:1501.07471 [hep-th]].
  %%CITATION = doi:10.1016/j.physletb.2015.09.013;%%
  %3 citations counted in INSPIRE as of 03 Apr 2019

%\cite{Fulling:2018lez}
\bibitem{Fulling:2018lez}
  S.~A.~Fulling and J.~H.~Wilson,
  ``The Equivalence Principle at Work in Radiation from Unaccelerated Atoms and Mirrors,''
  Phys.\ Scripta {\bf 94}, no. 1, 014004 (2019)
  doi:10.1088/1402-4896/aaecaa
  [arXiv:1805.01013 [quant-ph]].
  %%CITATION = doi:10.1088/1402-4896/aaecaa;%%
  %4 citations counted in INSPIRE as of 08 Jun 2019



%\cite{Zubarev}
\bibitem{Zubarev}  
D. N. Zubarev, A. V. Prozorkevich, S. A. Smolyanskii, "Derivation of nonlinear generalized equations of quantum relativistic hydrodynamics", TMF, 40:3 (1979), 394-407; Theoret. and Math. Phys., 40:3 (1979), 821-831.

%\cite{Buzzegoli:2017cqy}
\bibitem{Buzzegoli:2017cqy}
  M.~Buzzegoli, E.~Grossi and F.~Becattini,
  ``General equilibrium second-order hydrodynamic coefficients for free quantum fields,''
  JHEP {\bf 1710} (2017) 091
  doi:10.1007/JHEP10(2017)091
  [arXiv:1704.02808 [hep-th]].
  %%CITATION = doi:10.1007/JHEP10(2017)091;%%
  %2 citations counted in INSPIRE as of 01 Nov 2017

%\cite{Buzzegoli:2018wpy}
\bibitem{Buzzegoli:2018wpy} 
  M.~Buzzegoli and F.~Becattini,
  ``General thermodynamic equilibrium with axial chemical potential for the free Dirac field,''
  JHEP {\bf 1812}, 002 (2018)
  doi:10.1007/JHEP12(2018)002
  [arXiv:1807.02071 [hep-th]].
  %%CITATION = doi:10.1007/JHEP12(2018)002;%%
  %3 citations counted in INSPIRE as of 13 Mar 2019 

%\cite{Becattini:2015nva}
\bibitem{Becattini:2015nva} 
  F.~Becattini and E.~Grossi,
  ``Quantum corrections to the stress-energy tensor in thermodynamic equilibrium with acceleration,''
  Phys.\ Rev.\ D {\bf 92}, 045037 (2015)
  doi:10.1103/PhysRevD.92.045037
  [arXiv:1505.07760 [gr-qc]].
  %%CITATION = doi:10.1103/PhysRevD.92.045037;%%
  %34 citations counted in INSPIRE as of 06 Nov 2019

%\cite{Prokhorov:2018bql}
\bibitem{Prokhorov:2018bql} 
  G.~Y.~Prokhorov, O.~V.~Teryaev and V.~I.~Zakharov,
  ``Effects of rotation and acceleration in the axial current: density operator vs Wigner function,''
  JHEP {\bf 1902}, 146 (2019)
  doi:10.1007/JHEP02(2019)146
  [arXiv:1807.03584 [hep-th]].
  %%CITATION = doi:10.1007/JHEP02(2019)146;%%
  %3 citations counted in INSPIRE as of 13 Mar 2019

%\cite{Kharzeev:2012ph}
\bibitem{Kharzeev:2012ph}
  D.~E.~Kharzeev, K.~Landsteiner, A.~Schmitt and H.~U.~Yee,
  ``'Strongly interacting matter in magnetic fields': an overview,''
  Lect.\ Notes Phys.\  {\bf 871} (2013) doi:10.1007/978-3-642-37305-3\_1
  [arXiv:1211.6245 [hep-ph]].
  %%CITATION = doi:10.1007/978-3-642-37305-3_1;%%
  %144 citations counted in INSPIRE as of 04 Jul 2017

%\cite{Son:2009tf}
\bibitem{Son:2009tf} 
  D.~T.~Son and P.~Surowka,
  ``Hydrodynamics with Triangle Anomalies,''
  Phys.\ Rev.\ Lett.\  {\bf 103}, 191601 (2009)
  doi:10.1103/PhysRevLett.103.191601
  [arXiv:0906.5044 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevLett.103.191601;%%
  %604 citations counted in INSPIRE as of 03 Jul 2019

%\cite{Sadofyev:2010is}
\bibitem{Sadofyev:2010is} 
  A.~V.~Sadofyev, V.~I.~Shevchenko and V.~I.~Zakharov,
  ``Notes on chiral hydrodynamics within effective theory approach,''
  Phys.\ Rev.\ D {\bf 83}, 105025 (2011)
  doi:10.1103/PhysRevD.83.105025
  [arXiv:1012.1958 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.83.105025;%%
  %92 citations counted in INSPIRE as of 03 Jul 2019

%\cite{Stone:2018zel}
\bibitem{Stone:2018zel} 
  M.~Stone and J.~Kim,
  ``Mixed Anomalies: Chiral Vortical Effect and the Sommerfeld Expansion,''
  Phys.\ Rev.\ D {\bf 98}, no. 2, 025012 (2018)
  doi:10.1103/PhysRevD.98.025012
  [arXiv:1804.08668 [cond-mat.mes-hall]].
  %%CITATION = doi:10.1103/PhysRevD.98.025012;%%
  %6 citations counted in INSPIRE as of 30 Nov 2018


%\cite{Landsteiner:2012kd}
\bibitem{Landsteiner:2012kd}
  K.~Landsteiner, E.~Megias and F.~Pena-Benitez,
  ``Anomalous Transport from Kubo Formulae,''
  Lect.\ Notes Phys.\  {\bf 871} (2013) 433
  doi:10.1007/978-3-642-37305-3$\_$17
  [arXiv:1207.5808 [hep-th]].
  %%CITATION = doi:10.1007/978-3-642-37305-3_17;%%
  %81 citations counted in INSPIRE as of 18 Oct 2017

%\cite{Becattini:2017ljh}
\bibitem{Becattini:2017ljh} 
  F.~Becattini,
  ``Thermodynamic equilibrium with acceleration and the Unruh effect,''
  Phys.\ Rev.\ D {\bf 97}, no. 8, 085013 (2018)
  doi:10.1103/PhysRevD.97.085013
  [arXiv:1712.08031 [gr-qc]].
  %%CITATION = doi:10.1103/PhysRevD.97.085013;%%
  %9 citations counted in INSPIRE as of 12 Mar 2019
 
  
%\cite{Prokhorov:2019cik}
\bibitem{Prokhorov:2019cik} 
  G.~Y.~Prokhorov, O.~V.~Teryaev and V.~I.~Zakharov,
  ``Unruh effect for fermions from the Zubarev density operator,''
  Phys.\ Rev.\ D {\bf 99}, no. 7, 071901 (2019)
  doi:10.1103/PhysRevD.99.071901
  [arXiv:1903.09697 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.99.071901;%%

%\cite{Prokhorov:2019hif}
\bibitem{Prokhorov:2019hif} 
  G.~Y.~Prokhorov, O.~V.~Teryaev and V.~I.~Zakharov,
  ``Instability at Unruh temperature as manifestation of singularity in complex momentum plane,''
  arXiv:1906.03529 [hep-th].
  %%CITATION = ARXIV:1906.03529;%%

%\cite{Dowker:1994fi}
\bibitem{Dowker:1994fi}
  J.~S.~Dowker,
  ``Remarks on geometric entropy,''
  Class.\ Quant.\ Grav.\  {\bf 11}, L55 (1994)
  doi:10.1088/0264-9381/11/4/001
  [hep-th/9401159].
  %%CITATION = doi:10.1088/0264-9381/11/4/001;%%
  %76 citations counted in INSPIRE as of 11 Oct 2019

%\cite{Frolov:1987dz}
\bibitem{Frolov:1987dz} 
  V.~P.~Frolov and E.~M.~Serebryanyi,
  ``Vacuum Polarization in the Gravitational Field of a Cosmic String,''
  Phys.\ Rev.\ D {\bf 35}, 3779 (1987).
  doi:10.1103/PhysRevD.35.3779
  %%CITATION = doi:10.1103/PhysRevD.35.3779;%%
  %148 citations counted in INSPIRE as of 05 Nov 2019

%\cite{Becattini:2012tc}
\bibitem{Becattini:2012tc} 
  F.~Becattini,
  ``Covariant statistical mechanics and the stress-energy tensor,''
  Phys.\ Rev.\ Lett.\  {\bf 108}, 244502 (2012)
  doi:10.1103/PhysRevLett.108.244502
  [arXiv:1201.5278 [gr-qc]].
  %%CITATION = doi:10.1103/PhysRevLett.108.244502;%%
  %20 citations counted in INSPIRE as of 01 Jun 2018
 
%\cite{DeGroot:1980dk}
\bibitem{DeGroot:1980dk}
  S.~R.~De Groot, W.~A.~Van Leeuwen and C.~G.~Van Weert,
  ``Relativistic Kinetic Theory. Principles and Applications,''
  Amsterdam, Netherlands: North-holland ( 1980) 417p
  %61 citations counted in INSPIRE as of 27 Dec 2017
 
%\cite{Laine:2016hma}
\bibitem{Laine:2016hma} 
  M.~Laine and A.~Vuorinen,
  ``Basics of Thermal Field Theory,''
  Lect.\ Notes Phys.\  {\bf 925}, pp.1 (2016)
  doi:10.1007/978-3-319-31933-9
  [arXiv:1701.01554 [hep-ph]].
  %%CITATION = doi:10.1007/978-3-319-31933-9;%%
  %19 citations counted in INSPIRE as of 01 Jun 2018 

%\cite{Florkowski:2018myy}
\bibitem{Florkowski:2018myy} 
  W.~Florkowski, E.~Speranza and F.~Becattini,
  ``Perfect-fluid hydrodynamics with constant acceleration along the stream lines and spin polarization,''
  Acta Phys.\ Polon.\ B {\bf 49}, 1409 (2018)
  doi:10.5506/APhysPolB.49.1409
  [arXiv:1803.11098 [nucl-th]].
  %%CITATION = doi:10.5506/APhysPolB.49.1409;%%
  %18 citations counted in INSPIRE as of 12 Mar 2019

\end{thebibliography}

\documentclass[prd,preprint,tightenlines,floatfix,showpacs,preprintnumbers,nofootinbib,eqsecnum]{revtex4}

 \usepackage[dvips,final]{graphicx}
  \usepackage{amssymb}
   \usepackage{amsmath}
    \usepackage{amsfonts}
     \usepackage{epsfig}
      \usepackage{bm} % bold math

\DeclareMathOperator*{\SumInt}{%
\mathchoice%
  {\ooalign{$\displaystyle\sum$\cr\hidewidth$\displaystyle\int$\hidewidth\cr}}
  {\ooalign{\raisebox{.14\height}{\scalebox{.7}{$\textstyle\sum$}}\cr\hidewidth$\textstyle\int$\hidewidth\cr}}
  {\ooalign{\raisebox{.2\height}{\scalebox{.6}{$\scriptstyle\sum$}}\cr$\scriptstyle\int$\cr}}
  {\ooalign{\raisebox{.2\height}{\scalebox{.6}{$\scriptstyle\sum$}}\cr$\scriptstyle\int$\cr}}
}
\usepackage{slashed}

\begin{document}

%\begin{flushright}
%\vskip1cm
%\end{flushright}

%=================================================================
\title{Calculation of acceleration effects using the Zubarev density operator}

\author{Georgy Y. Prokhorov$^1$}
\email{prokhorov@theor.jinr.ru}

\author{Oleg V. Teryaev$^{1,2,3}$}
\email{teryaev@jinr.ru}

\author{Valentin I. Zakharov$^{2,4,5}$}
\email{vzakharov@itep.ru}
\affiliation{
{$^1$\sl
%Bogoliubov Laboratory of Theoretical Physics,
Joint Institute for Nuclear Research, 141980 Dubna,  Russia \\
$^2$ \sl
Institute of Theoretical and Experimental Physics, NRC Kurchatov Institute,
117218 Moscow, Russia
 \\
$^3$ \sl
National Research Nuclear University MEPhI (Moscow Engineering Physics Institute), Kashirskoe Shosse 31, 115409 Moscow, Russia
 \\
$^4$ \sl
School of Biomedicine, Far Eastern Federal University, 690950 Vladivostok, Russia\\
$^5$ \sl
Moscow Institute of Physics and Technology, 141700 Dolgoprudny, Russia
}\vspace{1.5cm}
}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{abstract}
\vspace{0.5cm} The relativistic form of the Zubarev density operator can be used to study quantum effects associated with acceleration of the medium. In particular, it was recently shown that the calculation of perturbative corrections in acceleration based on the Zubarev density operator makes it possible to show the existence of Unruh effect. In this paper, we present the details of the calculation of quantum correlators arising in the fourth order of the perturbation theory needed to demonstrate the Unruh effect.  Expressions for the quantum corrections for massive fermions are also obtained.\end{abstract}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\maketitle

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%================
\section{Introduction}
\label{sec:intro}
%================

There are wonderful quantum-field effects associated with non-uniform motion of the medium. A well-known example of such an effect is the Unruh effect, according to which an accelerated observer perceives the Minkowski vacuum as a medium filled with particles with a temperature depending on the acceleration $ T_U = |a|/(2\pi) $ \cite{Unruh:1976db}. This temperature is called the Unruh temperature. This effect is similar to the Hawking effect, since it is also associated with the appearance of the event horizon in the accelerated system. The Unruh effect continues to be the focus of theorists \cite{Castorina:2012yg, Ohsaku:2004rv, Takeuchi:2015nga, Fulling:2018lez}.

In \cite{Zubarev} a universal fundamental statistical approach to describing the equilibrium thermodynamics of quantized fields was developed. This approach is based on the relativistic form of the Zubarev density operator. It has recently been shown that this approach allows to study in a regular way the effects of rotation and acceleration in a medium of relativistic particles \cite{Buzzegoli:2017cqy, Buzzegoli:2018wpy, Becattini:2015nva}.

Using the Zubarev operator method, various effects associated with the motion of the medium are shown. In particular, the well-known chiral vortical effect is shown and corrections to this effect are calculated \cite{Buzzegoli:2017cqy, Buzzegoli:2018wpy, Prokhorov:2018bql}. Since the chiral vortical effect is associated with the axial electromagnetic anomaly \cite{Kharzeev:2012ph, Son:2009tf, Sadofyev:2010is}, as well as with the gravitational anomaly \cite{Landsteiner:2012kd, Stone:2018zel}, it turns out that the approach with the Zubarev operator carries information about the most fundamental properties of matter.

A remarkable observation made recently is that the Unruh effect can also be obtained from the Zubarev density operator \cite{Prokhorov:2019cik, Becattini:2017ljh}. This means, that in the Zubarev approach, nontrivial gravitational effects, associated with the occurrence of an event horizon, and the changes in vacuum properties depending on the reference system, are also reproduced.
This observation seems even more surprising because the corresponding calculation was carried out in ordinary flat Minkowski space, that is, by observing an accelerated medium from an inertial frame. Nevertheless, nontrivial physics associated with Unruh effect is reproduced.

Moreover, as discussed in \cite{Prokhorov:2019hif}, the Zubarev density operator  exactly reproduces quantum corrections that were derived in space of a cosmic string, characterized by a conical singularity \cite{Dowker:1994fi, Frolov:1987dz}. The existence of such exact duality means that the Zubarev operator of the accelerated medium leads to emergent conical geometry.

To justify the Unruh effect in \cite{Prokhorov:2019cik}, it was necessary to calculate a four-point correlator with boost operators. This calculation in \cite{Prokhorov:2019cik} was made for the massless Dirac field. In this paper, we give previously unpublished details of this calculation, and also derive expressions for the fourth-order corrections to the energy-momentum tensor at nonzero mass.

The paper has the following structure. Section \ref{sec:oper} introduces the basic concepts of the method of Zubarev density operator. An algorithm of constructing a perturbation theory in acceleration is also discussed. In the Section \ref{sec:coef} we describe in details the calculation of the corrections of the fourth-order in acceleration to the energy-momentum tensor. The interpretation associated with the Unruh effect is given in \ref{sec:disc}. In \ref{sec:concl} the conclusions are given. In the Appendix \ref{sec:m} the formulas for the coefficients at finite mass are presented.

The system of units $\hbar=c=k=1$ is used.

%================
\section{Perturbation theory in acceleration based on the Zubarev density operator}
\label{sec:oper}
%================

In this section, we introduce the basic concepts related to the density operator and describe how the acceleration perturbation theory can be constructed. In general, in this section we follow the paper \cite{Buzzegoli:2017cqy}. In \cite{Zubarev}, a relativistic form of the Zubarev density operator was obtained for a medium in a state of local thermodynamic equilibrium
\begin{eqnarray}
&&\hat{\rho}=\frac{1}{Z}\exp\Big\{
-\int_{\Sigma}d\Sigma_{\mu}[\hat{T}^{\mu\nu}(x)\beta_{\nu}(x)-\xi (x) \hat{j}^{\mu}(x)
]\Big\} \,,
\label{rho}
\end{eqnarray}
where integration over the 3-dimensional hypersurface  $\Sigma$ is performed. Here, $\beta_{\mu}=\frac{u_{\mu}}{T}$  is the inverse temperature 4-vector, $T$ is the proper temperature, $\xi =\frac{u}{T}$ is the ratio of the chemical potential in the comoving frame to temperature, $\hat{T}^{\mu\nu}$ and $\hat{j}^{\mu}$ are the energy-momentum tensor and current operators. The conditions of global thermodynamic equilibrium for a medium with rotation and acceleration, that is, conditions under which the density operator (\ref{rho}) becomes independent on the choice of the hypersurface $\Sigma$, over which the integration occurs, and thus acquires the properties of a density operator in a state of global thermodynamic equilibrium, have the form \cite{Becattini:2012tc, DeGroot:1980dk, Buzzegoli:2017cqy, Becattini:2015nva}
\begin{eqnarray}
&&\beta_{\mu}= b_{\mu}+\varpi_{\mu\nu}x_{\nu}\,,\quad b_{\mu}=\mathrm{const}\,,\quad\, \varpi_{\mu\nu}=\mathrm{const}\,,\nonumber \\
&&\varpi_{\mu\nu}= -\frac{1}{2}(\partial_{\mu}\beta_{\nu}-\partial_{\nu}\beta_{\mu})\,,\quad \xi=\text{const}\,,
\label{global}
\end{eqnarray}
where $\varpi_{\mu\nu}$ is the thermal vorticity tensor. Under the condition (\ref{global}) , the density operator (\ref{rho}) becomes the global equilibrium density operator \cite{Buzzegoli:2017cqy, Becattini:2015nva, Becattini:2017ljh}
\begin{eqnarray}
&&\hat{\rho}=\frac{1}{Z}\exp\Big\{-\beta_{\mu}(x)\hat{P}^{\mu}
+\frac{1}{2}\varpi_{\mu\nu}\hat{J}^{\mu\nu}_x+\xi  \hat{Q}
\Big\} \,,
\label{rho global}
\end{eqnarray}
where $\hat{P}$ is the 4-momentum operator, $\hat{Q}$ is the charge operator, and  $\hat{J}_x$ are the generators of the Lorentz transformations shifted to the point $x$
\begin{eqnarray}
&& \hat{J}^{\mu\nu}_x=
\int d\Sigma_{\lambda}\big[(y^{\mu}-x^{\mu})\hat{T}^{\lambda\nu}(y)-(y^{\nu}-x^{\nu})\hat{T}^{\lambda\mu}(y)\big] \,.
\label{J}
\end{eqnarray}

The technique of calculating the mean values of physical quantities based on (\ref{rho global}) was developed in \cite{Buzzegoli:2017cqy, Becattini:2015nva}, in which second-order corrections in the thermal vorticity tensor were calculated to various thermodynamic quantities for scalar and Dirac fields.

Note that the conditions (\ref{global}) also lead to a system of kinematic equations of motion, solving which, we can construct trajectories of motion. Particular cases of this solution are the rotation of the medium as a solid, as well as uniformly accelerated motion.

Following \cite{Buzzegoli:2017cqy}, we introduce the thermal acceleration vector $\alpha_{\mu}$ and the thermal vorticity pseudovector $w_{\mu}$
\begin{eqnarray}
\alpha_{\mu}= \varpi_{\mu\nu} u^{\nu},\quad w_{\mu}= -\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}u^{\nu}\varpi^{\alpha\beta}\,.
\label{alphaw}
\end{eqnarray}
Drawing a parallel with the electrodynamics,  $\alpha_{\mu}$ and $w_{\mu}$ can be called the "electrical" and "magnetic" components of the tensor $ \varpi $. The tensor $\varpi_{\mu\nu}$ can be
decomposed into these components as follows
\begin{eqnarray}
\varpi_{\mu\nu}= \epsilon_{\mu\nu\alpha\beta}w^{\alpha}u^{\beta}+\alpha_{\mu}u_{\nu}-\alpha_{\nu}u_{\mu}\,.
\label{decomp}
\end{eqnarray}
The meaning of the vectors $\alpha_{\mu}$ and $w_{\mu}$ becomes clear when considering the case of global thermodynamic equilibrium, in which they are proportional to the usual kinematic
4-acceleration $a_{\mu}$ and vorticity $\omega_{\mu}$
\begin{eqnarray}
\alpha^{\mu}={\varpi^{\mu}}_{\nu}u^{\nu}=u^{\nu}\partial_{\nu}\beta^{\mu}=\frac{1}{T}u^{\nu}\partial_{\nu}u^{\mu}=\frac{a^{\mu}}{T}\,,
\label{Add4}
\end{eqnarray}
and for thermal vorticity, we get
\begin{eqnarray}
w_{\mu}=-\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}u^{\nu}\varpi^{\alpha\beta}=
-\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}u^{\nu}\partial^{\beta}\beta^{\alpha}=
\frac{1}{2 T}\epsilon_{\mu\nu\alpha\beta}
u^{\nu}\partial^{\alpha}u^{\beta}=\frac{\omega_{\mu}}{T}\,.
\label{Add5}
\end{eqnarray}
In the rest frame, $a^{\mu}$ and $\omega^{\mu}$ are expressed in terms of 3-dimensional vectors
\begin{eqnarray}
a^{\mu}= (0,{\bf a}),\quad \omega^{\mu}= (0,{\bf w})\,,
\label{alphaw global rest}
\end{eqnarray}
where ${\bf a}$ and ${\bf w}$ are three-dimensional acceleration and angular velocity.

The density operator (\ref{rho global}) allows one to find corrections related to thermal vorticity in the framework of perturbation theory. To do this, it is necessary to expand (\ref{rho global}) in a series taking into account the fact that we are constructing a perturbation theory with noncommuting operators. According to \cite{Buzzegoli:2017cqy} we have
\begin{eqnarray}
\langle\hat{O}(x)\rangle=
\langle\hat{O}(0)\rangle_{\beta(x)}+
\sum_{N=1}^{\infty}\frac{\varpi^N }{2^N N!|\beta|^N}\int_{0}^{|\beta|}d\tau_1d\tau_2...d\tau_N
\langle T_{\tau} 
\hat{J}_{-i\tau_1 u}...
\hat{J}_{-i\tau_N u}
\hat{O}(0) 
\rangle_{\beta(x),c}
\,,
\label{On}
\end{eqnarray}
where it is assumed that each of the thermal vorticity tensors is contracted with the tensor $ \hat{J} $ so that  $ \varpi_{\mu\nu} \hat{J}^{\mu\nu}$. Eq.~(\ref{On}) includes only connected correlators, all disconnected correlators are reduced due to the contribution of the denominator $1/Z$ to (\ref{rho global}). This fact is reflected in the subscript $ c $, the subscript $ \beta(x) $ means that the mean values are taken at $\varpi = 0 $, that is, averaging is performed over a grand canonical distribution. The $T_{\tau}$ operator orders in imaginary time $\tau$, and $|\beta|=\sqrt{\beta^{\mu}\beta_{\mu}}=\frac{1}{T}$.

It is convenient to introduce the boost operator $\hat{K}$ and the angular momentum operator $\hat{J}$
\begin{eqnarray}
&& \hat{J}^{\mu\nu}=u^{\mu}\hat{K}^{\nu}-u^{\nu}\hat{K}^{\mu}-
\epsilon^{\mu\nu\rho\sigma}u_{\rho}\hat{J}_{\sigma} \,.
\label{Jdec}
\end{eqnarray}
From (\ref{decomp}) and (\ref{Jdec}) it follows that scalar products with vorticity tensor in (\ref{rho global}) and (\ref{On}) decompose into terms with boost and angular momentum
\begin{eqnarray}
\varpi_{\mu\nu} \hat{J}^{\mu\nu}_x=-2\alpha_{\mu}\hat{K}^{\mu}_x -2w_{\mu}\hat{J}^{\mu}_x \,.
\label{w}
\end{eqnarray}
Further, we will consider uniformly accelerated media without vorticity and chemical potential, therefore (\ref{rho global}) transforms to the density operator of the form
\begin{eqnarray}
\hat{\rho}=\frac{1}{Z}\exp\Big\{-\beta_{\mu}\hat{P}^{\mu}
-\alpha_{\mu} \hat{K}^{\mu}_x
\Big\} \,,
\label{operk}
\end{eqnarray}
and the perturbation theory in (\ref{On}), takes the form of the series in acceleration
\begin{eqnarray}
\langle\hat{O}(x)\rangle=
\langle\hat{O}(0)\rangle_{\beta(x)}+
\sum_{N=1}^{\infty}\frac{(-1)^N a^N }{ N!}\int_{0}^{|\beta|}d\tau_1 d\tau_2... d\tau_N
\langle T_{\tau} 
\hat{K}_{-i\tau_1 u}...
\hat{K}_{-i\tau_N u}
\hat{O}(0) 
\rangle_{\beta(x),c}
\,,
\label{On1}
\end{eqnarray}

%================
\section{Calculation of fourth-order coefficients in acceleration}
\label{sec:coef}
%================

The second-order coefficients in acceleration in the energy-momentum tensor of the Dirac field were calculated in \cite{Buzzegoli:2017cqy, Buzzegoli:2018wpy}. In this section, we present the details of calculation of the fourth-order coefficient.

The operator form of the energy-momentum tensor of the massless Dirac fields is well known. We will use the symmetrised Belinfante energy-momentum tensor
\begin{eqnarray}
\hat{T}^{\mu\nu}=\frac{i}{4}\left(
\overline{\Psi}\gamma^{\mu}\partial^{\nu}\Psi-
\partial^{\nu}\overline{\Psi}\gamma^{\mu}\Psi+
\overline{\Psi}\gamma^{\nu}\partial^{\mu}\Psi-
\partial^{\mu}\overline{\Psi}\gamma^{\nu}\Psi
\right)\,.
\label{Tmin}
\end{eqnarray}
As follows from (\ref{On1}), the calculation of the necessary correlators is performed in imaginary time - a time shift is made along the imaginary axis. Thus, it is necessary to pass to the Euclidean formalism in imaginary time. The Euclidean version of the energy-momentum tensor (\ref{Tmin}) has the form
\begin{eqnarray}
\hat{T}^{\mu\nu}=\frac{i^{\delta_{0\mu}+\delta_{0\nu}}}{4}\left(
\overline{\Psi}\tilde{\gamma}^{\mu}\partial^{\nu}\Psi-
\partial^{\nu}\overline{\Psi}\tilde{\gamma}^{\mu}\Psi+
\overline{\Psi}\tilde{\gamma}^{\nu}\partial^{\mu}\Psi-
\partial^{\mu}\overline{\Psi}\tilde{\gamma}^{\nu}\Psi
\right)\,,
\label{Teucl}
\end{eqnarray}
where $ \tilde{\gamma} $ are the Euclidean Dirac matrices
\begin{eqnarray}
\tilde{\gamma}_{\mu}=i^{1-\delta_{0\mu}}{\gamma}_{\mu}\,,\quad
\tilde{\gamma}^{\mu}=i^{1-\delta_{0\mu}}{\gamma}^{\mu} \,,\quad \{\tilde{\gamma}_{\mu}
\tilde{\gamma}_{\nu}\}=2\delta_{\mu\nu}\,,
\label{geuc}
\end{eqnarray}
and derivatives are also taken in Euclidean space-time, so that
\begin{eqnarray}
\tilde{\partial}_{\mu}=(-i)^{\delta_{0\mu}}{\partial}_{\mu}\,,
\end{eqnarray}
however, we will omit the tilde sign for derivatives. Consider the mean value of the energy-momentum tensor in the fourth order of the perturbation theory in acceleration using (\ref{On1})
\begin{eqnarray}
\langle\hat{T}^{\mu\nu}(x)\rangle &=& \langle\hat{T}^{\mu\nu}(0)\rangle_{\beta(x)}+
\frac{a_{\rho} a_{\sigma}}{2}
\int_{0}^{|\beta|}d\tau_1 d\tau_2
\langle T_{\tau} \hat{K}^{\rho}_{-i\tau_1 u}\hat{K}^{\sigma}_{-i\tau_2 u}\hat{T}^{\mu\nu}(0)\rangle_{\beta(x),c}
+ \label{T4} \\
&& +\frac{8 a_{\rho} a_{\sigma} a_{\gamma} a_{\eta}}{4!}
\int_0^{|\beta|}d\tau_1 d\tau_2 d\tau_3 d\tau_4
\langle T_{\tau} 
\hat{K}^{\rho}_{-i\tau_1 u}
\hat{K}^{\sigma}_{-i\tau_2 u}
\hat{K}^{\gamma}_{-i\tau_3 u}
\hat{K}^{\eta}_{-i\tau_4 u}
\hat{T}^{\mu\nu}(0)\rangle_{\beta(x),c}
+ \mathcal{O}(a^6)\,.  \nonumber
\end{eqnarray}

Symmetry and parity considerations fix the form of the energy-momentum tensor in the fourth order of perturbation theory
\begin{eqnarray}
\langle \hat{T}^{\mu\nu}\rangle &=& (\rho_0 - A_1 T^2 a^2   +A_2 a^4 )u^{\mu}u^{\nu}-
(p_0-A_3 T^2 a^2  + A_4 a^4)\Delta^{\mu\nu} \nonumber \\
&&+(A_5 T^2- A_6 a^2)a^{\mu}a^{\nu}+\mathcal{O}(a^6) \qquad
\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}\,,
\label{Ta4}
\end{eqnarray}
where $ a^2=a_{\mu}a^{\mu} $. As already mentioned, 2-order coefficients were calculated earlier in \cite{Buzzegoli:2017cqy, Buzzegoli:2018wpy}. Our goal is to calculate coefficients of the 4th order $ A_2, A_4, A_6 $. Comparing (\ref{T4}) and (\ref{Ta4}), we obtain
\begin{eqnarray}
A_2 a^4 u^{\mu}u^{\nu}-
A_4 a^4\Delta^{\mu\nu}- A_6 a^2 a^{\mu}a^{\nu} &=&
\frac{a_{\rho}a_{\sigma}a_{\gamma}a_{\eta}}{4!}
\int_0^{|\beta|}d\tau_1 d\tau_2 d\tau_3 d\tau_4  \nonumber \\
&&\times\langle T_{\tau} 
\hat{K}^{\rho}_{-i\tau_1 u}
\hat{K}^{\sigma}_{-i\tau_2 u}
\hat{K}^{\gamma}_{-i\tau_3 u}
\hat{K}^{\eta}_{-i\tau_4 u}
\hat{T}^{\mu\nu}(0)\rangle_{\beta(x),c}\,.
\label{corr}
\end{eqnarray}
The coefficients $ A_2, A_4, A_6 $ are Lorentz invariants, and the relation (\ref{corr})  is valid for any choice of the vectors $ u_{\mu}, a_{\mu} $. Therefore, to determine the coefficient, we can choose the vectors $ u_{\mu}, a_{\mu} $ in any form convenient for us. To determine $ A_2 $, we choose $ a^{\mu}=(0,0,0,|a|) $ and $ u^{\mu}=(1,0,0,0) $ and consider the components  $ \mu=0, \nu=0 $, to determine $ A_4 $ we choose $ a^{\mu}=(0,0,|a|,0) $ and $ u^{\mu}=(1,0,0,0) $ and consider the components $ \mu=3, \nu=3 $, to determine $ A_6 $ we choose $ a^{\mu}=(0,0,0,|a|) $ and $ u^{\mu}=(1,0,0,0) $ and consider the components $ \mu=3, \nu=3 $. As a result, we obtain
\begin{eqnarray}
A_2 &=& \frac{1}{4!}
\int_0^{|\beta|}d\tau_1 d\tau_2 d\tau_3 d\tau_4 \langle T_{\tau} 
\hat{K}^{3}_{-i\tau_1 u}
\hat{K}^{3}_{-i\tau_2 u}
\hat{K}^{3}_{-i\tau_3 u}
\hat{K}^{3}_{-i\tau_4 u}
\hat{T}^{00}(0)\rangle_{\beta(x),c}\,,  \nonumber \\
A_4 &=& \frac{1}{4!}
\int_0^{|\beta|}d\tau_1 d\tau_2 d\tau_3 d\tau_4 \langle T_{\tau} 
\hat{K}^{2}_{-i\tau_1 u}
\hat{K}^{2}_{-i\tau_2 u}
\hat{K}^{2}_{-i\tau_3 u}
\hat{K}^{2}_{-i\tau_4 u}
\hat{T}^{33}(0)\rangle_{\beta(x),c}\,,  \nonumber \\
A_6 &=&-A_4+ \frac{1}{4!}
\int_0^{|\beta|}d\tau_1 d\tau_2 d\tau_3 d\tau_4 \langle T_{\tau} 
\hat{K}^{3}_{-i\tau_1 u}
\hat{K}^{3}_{-i\tau_2 u}
\hat{K}^{3}_{-i\tau_3 u}
\hat{K}^{3}_{-i\tau_4 u}
\hat{T}^{33}(0)\rangle_{\beta(x),c}\,.
\label{afirst}
\end{eqnarray}
We now use the representation of the boost operator through the energy-momentum tensor. According to (\ref{J}) and (\ref{Jdec}) we have
\begin{eqnarray}
\hat{K}^3_{-i\tau u} &=&  \hat{J}^{03}_{-i\tau u}
=
\int d^3 x(-1) x^3 \hat{T}^{00}(\tau,{\bf x})\,, \nonumber \\
\hat{K}^2_{-i\tau u} &=&  \hat{J}^{02}_{-i\tau u}
=
\int d^3 x (-1) x^2 \hat{T}^{00}(\tau,{\bf x})\,, 
\label{J3K3K1}
\end{eqnarray}
Substituting (\ref{J3K3K1}) into (\ref{afirst}), we come to the need of calculating quantities of the form
\begin{eqnarray} 
&&C^{\alpha_1\alpha_2|\alpha_3\alpha_4
|\alpha_5\alpha_6|\alpha_7\alpha_8|\alpha_9\alpha_{10}|
ijkl} =\int_0^{|\beta|} d\tau_xd\tau_yd\tau_zd\tau_f d^3x d^3y d^3z d^3f \nonumber \\
&&\times x^{i} y^{j} z^{k} f^{l}\langle
T_{\tau} \hat{T}^{\alpha_1\alpha_2}(\tau_x,{\bf x})\hat{T}^{\alpha_3\alpha_4}(\tau_y,{\bf y})\hat{T}^{\alpha_5\alpha_6}(\tau_z,{\bf z})
\hat{T}^{\alpha_7\alpha_8}(\tau_f,{\bf f})\hat{T}^{\alpha_9\alpha_{10}}(0)
\rangle_{\beta(x),c}\,.
\label{C}
\end{eqnarray}
In particular, from  (\ref{afirst}), we have
\begin{eqnarray} 
A_2 = \frac{1}{4!}C^{00|00|00|00|00|3333},\,A_4 = \frac{1}{4!}C^{00|00|00|00|33|2222},\,A_6 =-A_4+ \frac{1}{4!}C^{00|00|00|00|33|3333}
.
\label{AC}
\end{eqnarray}
Next, we will focus on calculating the coefficient in energy $ A_2 $, the remaining coefficients can be calculated by analogy.

We represent the energy-momentum tensor (\ref{Teucl}) in a split form
\begin{eqnarray}
\hat{T}^{\alpha\beta}(X) &=& \lim_{\scriptscriptstyle X_1,X_2\to X} \mathcal{D}^{\alpha\beta}_{a b}(\partial_{X_1},\partial_{X_2})\bar{\Psi}_{a}(X_1)\Psi_{b}(X_2)\,,\nonumber \\ 
\mathcal{D}^{\alpha\beta}_{a b}(\partial_{X_1},\partial_{X_2})&=&\frac{i^{\delta_{0\alpha}+\delta_{0\beta}}}{4}
[\tilde{\gamma}_{a b}^{\alpha}(\partial_{X_2}-\partial_{X_1})^{\beta}+\tilde{\gamma}_{a b}^{\beta}(\partial_{X_2}-\partial_{X_1})^{\alpha}]\,,
\label{Tsplit}
\end{eqnarray}
and substitute it in (\ref{AC}). As a result, we get for the corresponding correlator
\begin{eqnarray}
&&\langle
T_{\tau}\hat{T}^{00}(X)\hat{T}^{00}(Y)\hat{T}^{00}(Z)\hat{T}^{00}(F)\hat{T}^{00}(0)\rangle_{\beta(x),c}=\lim_{
\def\arraystretch{0.5}\begin{array}{ll}
{\scriptscriptstyle X_1,X_2\to X}\vspace{0.1mm}\\
{\scriptscriptstyle Y_1,Y_2\to Y}\\
{\scriptscriptstyle Z_1,Z_2\to Z}\\
{\scriptscriptstyle F_1,F_2\to F}\\
{\scriptscriptstyle H_1,H_2\to H=0}
\end{array}} \mathcal{D}^{00}_{a_1 a_2}(\partial_{X_1},\partial_{X_2}) \nonumber \\
&&\mathcal{D}^{00}_{a_3 a_4}(\partial_{Y_1},\partial_{Y_2})\mathcal{D}^{00}_{a_5 a_6}(\partial_{Z_1},\partial_{Z_2})
\mathcal{D}^{00}_{a_7 a_8}(\partial_{F_1},\partial_{F_2})
\mathcal{D}^{00}_{a_9 a_{10}}(\partial_{H_1},\partial_{H_2})
\langle T_{\tau} \overline{\Psi}_{a_1}(X_1)\Psi_{a_2}(X_2)\nonumber \\ 
&&\times \overline{\Psi}_{a_3}(Y_1)\Psi_{a_4}(Y_2) \overline{\Psi}_{a_5}(Z_1)\Psi_{a_6}(Z_2)\overline{\Psi}_{a_7}(F_1)\Psi_{a_8}(F_2)
\overline{\Psi}_{a_9}(H_1)\Psi_{a_{10}}(H_2)\rangle_{\beta(x),c}\,.
\label{corrsplit}
\end{eqnarray}
When calculating the correlator with 10 Dirac fields of the form (\ref{corrsplit}) , it is necessary to use an analogue of Wick theorem for field theory at finite temperatures. Then the five-point correlator in (\ref{corrsplit}) leads to the product of mean values of quadratic combinations of Dirac fields, that is, thermal propagators. For short, we denote  $\Psi_{a_n}\to n$, and $ \overline{\Psi}_{a_n}\to \bar{n}$ and omit  $T_{\tau}$ and the index $\beta(x)$. Then, after extraction on the connected part in (\ref{corrsplit}) according to Wick theorem, we obtain 24 terms
\begin{eqnarray}
&&\langle T_{\tau} \overline{\Psi}_{a_1}(X_1)\Psi_{a_2}(X_2)\overline{\Psi}_{a_3}(Y_1)\Psi_{a_4}(Y_2)\overline{\Psi}_{a_5}(Z_1)\Psi_{a_6}(Z_2)\overline{\Psi}_{a_7}(F_1)\Psi_{a_8}(F_2) \nonumber \\
&&\overline{\Psi}_{a_9}(H_1)\Psi_{a_{10}}(H_2)\rangle_{\beta(x),c}
=\langle\bar{1}2\bar{3}4\bar{5}6\bar{7}8\bar{9}10\rangle = 
- 
 \langle \bar{1}  4\rangle  \langle 2    \bar{9}\rangle  \langle  \bar{3}   6\rangle  \langle  \bar{5}   8\rangle  \langle  \bar{7}   10\rangle\nonumber \\
&& 
+
 \langle \bar{1}  4\rangle  \langle 2   \bar{7}\rangle  \langle \bar{3}  6\rangle  \langle \bar{5}  10\rangle  \langle 8   \bar{9}\rangle  
+
 \langle \bar{1}  4\rangle  \langle 2   \bar{9}\rangle  \langle \bar{3} 8\rangle  \langle \bar{5}  10\rangle  \langle 6  \bar{7}\rangle 
+
 \langle \bar{1}  4\rangle  \langle 2   \bar{5}\rangle  \langle \bar{3} 8\rangle  \langle 6   \bar{9}\rangle  \langle \bar{7}  10\rangle\nonumber \\
&& 
+
 \langle \bar{1}  4\rangle  \langle 2   \bar{7}\rangle  \langle \bar{3}  10\rangle  \langle \bar{5}  8\rangle  \langle 6  \bar{9}\rangle 
 -
 \langle \bar{1}  4\rangle  \langle 2    \bar{5}\rangle  \langle  \bar{3}   10\rangle  \langle 6    \bar{7}\rangle  \langle 8    \bar{9}\rangle  
+
 \langle \bar{1}  6\rangle  \langle 2   \bar{9}\rangle  \langle \bar{3}  8\rangle  \langle 4   \bar{5}\rangle  \langle \bar{7} 10\rangle \nonumber \\
&& 
- 
 \langle  \bar{1}   6\rangle  \langle 2    \bar{7}\rangle  \langle  \bar{3}   10\rangle  \langle 4    \bar{5}\rangle  \langle 8    \bar{9}\rangle  
+
 \langle \bar{1}  6\rangle  \langle 2   \bar{9}\rangle  \langle \bar{3}  10\rangle  \langle 4   \bar{7}\rangle  \langle \bar{5}  8\rangle  
+
 \langle \bar{1}  6\rangle  \langle 2   \bar{3}\rangle  \langle 4  \bar{9}\rangle  \langle \bar{5}  8\rangle  \langle \bar{7}  10\rangle  \nonumber \\
&&
+
 \langle \bar{1}  6\rangle  \langle 2  \bar{7}\rangle  \langle \bar{3}  8\rangle  \langle 4   \bar{9}\rangle  \langle \bar{5}  10\rangle  
- 
 \langle  \bar{1}   6\rangle  \langle 2    \bar{3}\rangle  \langle 4    \bar{7}\rangle  \langle  \bar{5}   10\rangle  \langle 8    \bar{9}\rangle  
+
 \langle \bar{1}  8\rangle  \langle 2  \bar{9}\rangle  \langle \bar{3}  6\rangle  \langle 4   \bar{7}\rangle  \langle \bar{5} 10\rangle  \nonumber \\
&&
- 
 \langle  \bar{1}   8\rangle  \langle 2    \bar{5}\rangle  \langle  \bar{3}   10\rangle  \langle 4    \bar{7}\rangle  \langle 6    \bar{9}\rangle  
- 
 \langle  \bar{1}   8\rangle  \langle 2    \bar{9}\rangle  \langle  \bar{3}   10\rangle  \langle 4    \bar{5}\rangle  \langle 6    \bar{7}\rangle  
- 
 \langle  \bar{1}   8\rangle  \langle 2    \bar{3}\rangle  \langle 4    \bar{9}\rangle  \langle  \bar{5}   10\rangle  \langle 6    \bar{7}\rangle \nonumber \\
&& 
+
 \langle \bar{1}  8\rangle  \langle 2  \bar{5}\rangle  \langle \bar{3}  6\rangle  \langle 4   \bar{9}\rangle  \langle \bar{7} 10\rangle  
- 
 \langle  \bar{1}   8\rangle  \langle 2    \bar{3}\rangle  \langle 4    \bar{5}\rangle  \langle 6    \bar{9}\rangle  \langle  \bar{7}   10\rangle  
+
 \langle \bar{1}  10\rangle  \langle 2  \bar{7}\rangle  \langle \bar{3}  6\rangle  \langle 4   \bar{9}\rangle  \langle \bar{5} 8\rangle  \nonumber \\
&&
- 
 \langle  \bar{1}   10\rangle  \langle 2    \bar{5}\rangle  \langle  \bar{3}   8\rangle  \langle 4    \bar{9}\rangle  \langle 6    \bar{7}\rangle  
- 
 \langle  \bar{1}   10\rangle  \langle 2    \bar{7}\rangle  \langle  \bar{3}   8\rangle  \langle 4    \bar{5}\rangle  \langle 6    \bar{9}\rangle  
- 
 \langle  \bar{1}   10\rangle  \langle 2    \bar{3}\rangle  \langle 4    \bar{7}\rangle  \langle  \bar{5}   8\rangle  \langle 6    \bar{9}\rangle \nonumber \\
&& 
- 
 \langle  \bar{1}   10\rangle  \langle 2    \bar{5}\rangle  \langle  \bar{3}   6\rangle  \langle 4    \bar{7}\rangle  \langle 8    \bar{9}\rangle  
+
 \langle \bar{1}  10\rangle  \langle 2  \bar{3}\rangle  \langle 4   \bar{5}\rangle  \langle 6   \bar{7}\rangle  \langle 8   \bar{9}\rangle
\,,
\label{wik}
\end{eqnarray}
where signs correspond to the number of permutations of anti-commuting fields. Thermal propagators have a standard form \cite{Buzzegoli:2017cqy, Laine:2016hma}
\begin{eqnarray}
G_{a_1a_2}(X_1,X_2)&=& \langle T_{\tau}\Psi_{a_1}(X_1)\overline{\Psi}_{a_2}(X_2)\rangle_{\beta(x)}=\SumInt_{P}e^{iP^+(X_1-X_2)}(-iP_{\mu}^+\tilde{\gamma}_{\mu}+m)_{a_1a_2}\Delta(P^+)\,, \nonumber \\
\bar{G}_{a_1a_2}(X_1,X_2)&=& \langle T_{\tau}\overline{\Psi}_{a_1}(X_1)\Psi_{a_2}(X_2)\rangle_{\beta(x)}=-\langle T_{\tau}\Psi_{a_2}(X_2)\overline{\Psi}_{a_1}(X_1)\rangle_{\beta(x)} \nonumber \\
&=& -\SumInt_{P}e^{iP^-(X_1-X_2)}(iP_{\mu}^-\tilde{\gamma}_{\mu}+m)_{a_2a_1}\Delta(P^-)\,,
\label{prop}
\end{eqnarray}
where integration over the three-dimensional components of the momentum and summation over the Matsubara frequencies of fermion field appear. The following notations are used in (\ref{prop}): $P^\pm=(p^{\pm}_n,{\bf p})$, $ p^{\pm}_n=\pi(2n+1)/|\beta|\pm i\mu $, $ n=0,\pm 1,\pm 2, ... $, $X=(\tau,{\bf x})$, $\SumInt_{P}=\frac{1}{|\beta|}\sum_{n=-\infty}^{\infty}\int\frac{d^3p}{(2\pi)^3}$, and $\Delta(P)=\frac{1}{P^2+m^2}$, where the square is taken with the Euclidean metric, as also in $P_{\mu}^{\pm}\tilde{\gamma}_{\mu}=\slashed{P}^{\pm}$ (unlike $P^+(X_1-X_2)$ according to \cite{Laine:2016hma}). Since we consider massless field at zero chemical potential, the mass and chemical potential must be set equal to zero $ m=0, \mu=0 $. Nevertheless, we retain the notation  $ P^\pm $, bearing in mind the possibility of generalization to the case with nonzero chemical potential in the future.

Next, substitute (\ref{prop}) in (\ref{wik}). We will describe the calculations for the first term in (\ref{wik}), while all other terms can be calculated by analogy
\begin{eqnarray}
&&-\lim_{
\def\arraystretch{0.5}\begin{array}{ll}
{\scriptscriptstyle X_1,X_2\to X}\vspace{0.1mm}\\
{\scriptscriptstyle Y_1,Y_2\to Y}\\
{\scriptscriptstyle Z_1,Z_2\to Z}\\
{\scriptscriptstyle F_1,F_2\to F}\\
{\scriptscriptstyle H_1,H_2\to H=0}
\end{array}} 
\mathcal{D}^{00}_{a_1 a_2}(\partial_{X_1},\partial_{X_2}) \mathcal{D}^{00}_{a_3 a_4}(\partial_{Y_1},\partial_{Y_2})
\mathcal{D}^{00}_{a_5 a_6}(\partial_{Z_1},\partial_{Z_2})
\mathcal{D}^{00}_{a_7 a_8}(\partial_{F_1},\partial_{F_2})
\mathcal{D}^{00}_{a_9 a_{10}}(\partial_{H_1},\partial_{H_2})\nonumber \\
&&\times
\bar{G}_{a_1a_4}(X_1,Y_2) 
G_{a_2a_9}(X_2,H_1)
\bar{G}_{a_3a_6}(Y_1,Z_2)
\bar{G}_{a_5a_8}(Z_1,F_2)
\bar{G}_{a_7a_{10}}(F_1,H_2)
= \nonumber \\
&&-\SumInt_{\{P,Q,K,R,L\}}e^{-i{\bf p}({\bf x}-{\bf y})-i{\bf q}{\bf x}-i{\bf k}({\bf y}-{\bf z})-i{\bf r}({\bf z}-{\bf f})-i{\bf l}{\bf f}} e^{ip^{-}_n(\tau_x -\tau_y)+iq^{+}_n\tau_x+ik^{-}_n (\tau_y-\tau_z) +ir^{-}_n(\tau_z-\tau_f)+il^{-}_n\tau_f}\nonumber \\
&&\times \Delta(P^{-})\Delta(Q^{+})\Delta(K^{-}) \Delta(R^{-})\Delta(L^{-})\nonumber \\
&&\times\mathrm{tr}\Big[
(-i\slashed{L}^{-})
\mathcal{D}^{00}(iL^{-},-iR^{-})
(-i\slashed{R}^{-})
\mathcal{D}^{00}(iR^{-},-iK^{-})
(-i\slashed{K}^{-}) 
\nonumber \\
&&\times
\mathcal{D}^{00}(iK^{-},-iP^{-})
(-i\slashed{P}^{-})
\mathcal{D}^{00}(iP^{-},iQ^{+})
(-i\slashed{Q}^{+})
\mathcal{D}^{00}(-iQ^{+},-iL^{-})
\Big]\,,
\label{inter1}
\end{eqnarray}
where it was necessary to arrange all the matrices under the trace in accordance with the order of the spinor indices. To calculate (\ref{inter1}), it is necessary to find a trace of the form
\begin{eqnarray}
\mathrm{tr}\Big[
 \slashed{P_1}
\mathcal{D}^{00}(P_2,P_3)
 \slashed{P_4}
\mathcal{D}^{00}(P_5 ,P_6 )
 \slashed{P_7}
\mathcal{D}^{00}(P_8 ,P_9 )
 \slashed{P_{10}}
\mathcal{D}^{00}(P_{11} ,P_{12} )
 \slashed{P_{13}}
\mathcal{D}^{00}(P_{14} ,P_{15} )
\Big]\,.
\label{typtr}
\end{eqnarray}
The subsequent calculations are more convenient to carry out using special software applications. Calculation (\ref{typtr}) requires finding the trace of 10 Euclidean Dirac matrices
\begin{eqnarray}
\mathrm{tr}\Big[
\tilde{\gamma}^{\alpha_1}
\tilde{\gamma}^{\alpha_2}
\tilde{\gamma}^{\alpha_3}
\tilde{\gamma}^{\alpha_4}
\tilde{\gamma}^{\alpha_5}
\tilde{\gamma}^{\alpha_6}
\tilde{\gamma}^{\alpha_7}
\tilde{\gamma}^{\alpha_8}
\tilde{\gamma}^{\alpha_9}
\tilde{\gamma}^{\alpha_{10}}
\Big]\,.
\end{eqnarray}
Using the definition (\ref{geuc}), this trace can be easily transformed to the trace of ordinary Dirac matrices, which can be found using standard methods. We denote the trace in (\ref{typtr}) as $ A(P_1,P_2,P_3,P_4,P_5,P_6,P_7,P_8,P_9,P_{10},P_{11},P_{12},P_{13},P_{14},P_{15}) $. Then (\ref{inter1}) will be presented in the form
\begin{eqnarray}
&&-\int \frac{d^3pd^3qd^3kd^3rd^3l}{(2\pi)^{15}} e^{-i{\bf p}({\bf x}-{\bf y})-i{\bf q}{\bf x}-i{\bf k}({\bf y}-{\bf z})-i{\bf r}({\bf z}-{\bf f})-i{\bf l}{\bf f}} \nonumber \\
&&
\times \sum_{p_n, q_n, k_n, r_n,l_n}\frac{1}{|\beta|^5} e^{ip^{-}_n(\tau_x -\tau_y)+iq^{+}_n\tau_x+ik^{-}_n (\tau_y-\tau_z) +ir^{-}_n(\tau_z-\tau_f)+il^{-}_n\tau_f}\nonumber \\
&&\times \Delta(P^{-})\Delta(Q^{+})\Delta(K^{-}) \Delta(R^{-})\Delta(L^{-})\nonumber \\
&&\times A(-iL^{-},iL^{-},-iR^{-},-iR^{-},iR^{-},-iK^{-},-iK^{-},iK^{-},-iP^{-},-iP^{-},iP^{-},iQ^{+},\nonumber \\
&&-iQ^{+},-iQ^{+},-iL^{-})\,.
\label{inter2}
\end{eqnarray}
Next, one needs to sum over the Matsubara frequencies in (\ref{inter2}) using the relation
\begin{eqnarray}
&&\frac{1}{|\beta|}\sum_{\omega_n}\frac{(\omega_n\pm i\mu)^k e^{i(\omega_n\pm i\mu)\tau}}{(\omega_n\pm i\mu)^2+E^2}=\frac{1}{2E}\sum_{s=\pm 1}(-isE)^k e^{\tau s E}[\theta(-s\tau)-n_F(E\pm s\mu)]\,,
\end{eqnarray}
where $ E=\sqrt{{\bf p}^2+m^2} $, $ n_F(E)=1/(1+e^{E/T}) $ is the Fermi distribution and $ \theta $ is the Heaviside theta function. Again, we can take $ m=0, \mu=0 $. As a result, we obtain
\begin{eqnarray}
&&-\int \frac{d^3pd^3qd^3kd^3rd^3l}{(2\pi)^{15}}e^{-i{\bf p}({\bf x}-{\bf y})-i{\bf q}{\bf x}-i{\bf k}({\bf y}-{\bf z})-i{\bf r}({\bf z}-{\bf f})-i{\bf l}{\bf f}} \nonumber \\
&&
\times \frac{1}{32 E_p E_q E_k E_r E_l}\sum_{s_1, s_2, s_3, s_4, s_5} e^{(\tau_x -\tau_y)s_1 E_p+\tau_x s_2 E_q+(\tau_y-\tau_z)s_3 E_k +(\tau_z-\tau_f)s_4 E_r+\tau_f s_5 E_l}\nonumber \\
&&\times A(-i\widetilde{L},i\widetilde{L},-i\widetilde{R},-i\widetilde{R},i\widetilde{R},-i\widetilde{K},-i\widetilde{K},i\widetilde{K},-i\widetilde{P},-i\widetilde{P},i\widetilde{P},i\widetilde{Q},-i\widetilde{Q},-i\widetilde{Q},-i\widetilde{L})\nonumber \\
&&\times\left(\theta\left[-s_1(\tau_x-\tau_y)\right]-n_F\left(E_p\right)\right)
\left(\theta\left[-s_2\right]-n_F\left(E_q\right)\right)\nonumber \\
&&\times 
\left( \theta\left[-s_3(\tau_y-\tau_z)\right]-n_F\left(E_k\right)\right)
\left( \theta\left[-s_4(\tau_z-\tau_f)\right]-n_F\left(E_r\right)\right)
\left( \theta\left[-s_5\right]-n_F\left(E_l\right)\right)
\,.
\label{inter3}
\end{eqnarray}
Here, following  \cite{Buzzegoli:2017cqy}, the notations  $\widetilde{P}=\widetilde{P}(s_1)=(-is_1 E_p,{\bf p})$, $\widetilde{Q}=\widetilde{Q}(s_2)$, ...  are introduced.
We return now to the formula for $ A_2 $ (\ref{AC}) with spatial integrals and calculate the contribution of the term (\ref{inter3}). This contribution has the form
\begin{eqnarray}
A_2&=&\int \frac{d\tau_x d\tau_y d\tau_z d\tau_f d^3 x d^3 y d^3 z d^3 f d^3pd^3qd^3kd^3r d^3l}{4!(2\pi)^{15}}
e^{-i{\bf p}({\bf x}-{\bf y})-i{\bf q}{\bf x}-i{\bf k}({\bf y}-{\bf z})-i{\bf r}({\bf z}-{\bf f})-i{\bf l}{\bf f}} \nonumber \\
&&\times x^{3}y^{3}z^{3}f^{3}D
+...\,,
\label{A2n}
\end{eqnarray}
where the ellipsis indicates the contribution of the remaining 23 terms from (\ref{wik}), and  $ D $ equals to
\begin{eqnarray}
D &=& - \frac{1}{32 E_p E_q E_k E_r E_l} \sum_{s_1, s_2, s_3, s_4, s_5} e^{(\tau_x -\tau_y)s_1 E_p+\tau_x s_2 E_q+(\tau_y-\tau_z)s_3 E_k +(\tau_z-\tau_f)s_4 E_r+\tau_f s_5 E_l} \label{D} \\
&&\times A(-i\widetilde{L},i\widetilde{L},-i\widetilde{R},-i\widetilde{R},i\widetilde{R},-i\widetilde{K},-i\widetilde{K},i\widetilde{K},-i\widetilde{P},-i\widetilde{P},i\widetilde{P},i\widetilde{Q},-i\widetilde{Q},-i\widetilde{Q},-i\widetilde{L})\nonumber \\
&&\times\left(\theta\left[-s_1(\tau_x-\tau_y)\right]-n_F\left(E_p\right)\right)
\left(\theta\left[-s_2\right]-n_F\left(E_q\right)\right) \nonumber \\
&&\times 
\left( \theta\left[-s_3(\tau_y-\tau_z)\right]-n_F\left(E_k\right)\right)
\left( \theta\left[-s_4(\tau_z-\tau_f)\right]-n_F\left(E_r\right)\right)
\left( \theta\left[-s_5\right]-n_F\left(E_l\right)\right)
\,. \nonumber
\end{eqnarray}
Next, one needs to rewrite the product of spatial coordinates in the integral through derivatives using the formula
\begin{eqnarray}
&&\int d^3p d^3q d^3k d^3r d^3x d^3y d^3z d^3f \,F({\bf p},{\bf q},{\bf k},{\bf r},{\bf l})e^{-i{\bf p}({\bf x}-{\bf y})-i{\bf q}{\bf x}-i{\bf k}({\bf y}-{\bf z})-i{\bf r}({\bf z}-{\bf f})-i{\bf l}{\bf f}} x^{3}y^{3}z^{3}f^{3} \nonumber \\
&&=(2\pi)^{12}\int d^3p \Big(-
\frac{\partial^3}{\partial q^{3}\partial l^{3}\partial p^{3}\partial r^{3}}
-\frac{\partial^3}{\partial q^{3}\partial l^{3}\partial p^{3}\partial l^{3}}
\nonumber \\
&&+\frac{\partial^3}{\partial q^{3}\partial l^{3}\partial r^{3}\partial q^{3}}
+\frac{\partial^3}{\partial q^{3}\partial q^{3}\partial l^{3}\partial l^{3}}
\Big)F({\bf p},{\bf q},{\bf k},{\bf r},{\bf l})\Bigg|\def\arraystretch{0.5}\begin{array}{ll}
{\scriptstyle {\bf l}={\bf p}}\\
{\scriptstyle {\bf r}={\bf p}}\\
{\scriptstyle {\bf k}={\bf p}}\\
{\scriptstyle {\bf q}=-{\bf p}}
\end{array}\,,
\label{dif rel}
\end{eqnarray}
resulting from integration by parts and properties of the delta function. After that  (\ref{A2n})  is converted to the form
\begin{eqnarray}\nonumber
A_2 &=& \frac{1}{4!(2\pi)^3} \int d\tau_x d\tau_y d\tau_z d\tau_f d^3 p \Big(-
\frac{\partial^3}{\partial q^{3}\partial l^{3}\partial p^{3}\partial r^{3}}
-\frac{\partial^3}{\partial q^{3}\partial l^{3}\partial p^{3}\partial l^{3}}
\nonumber \\
&&+\frac{\partial^3}{\partial q^{3}\partial l^{3}\partial r^{3}\partial q^{3}}
+\frac{\partial^3}{\partial q^{3}\partial q^{3}\partial l^{3}\partial l^{3}}
\Big)D({\bf p},{\bf q},{\bf k},{\bf r},{\bf l})\Bigg|\def\arraystretch{0.5}\begin{array}{ll}
{\scriptstyle {\bf l}={\bf p}}\\
{\scriptstyle {\bf r}={\bf p}}\\
{\scriptstyle {\bf k}={\bf p}}\\
{\scriptstyle {\bf q}=-{\bf p}}
\end{array}+...\,.
\label{A2n1}
\end{eqnarray}

Now, it remains to integrate over the imaginary time and also over the last momentum, which can be done directly in spherical coordinates $ d^3p=|{\bf p}|^2 d|{\bf p}| \sin(\theta) d\theta d\phi $. The sequence of actions in this case, from the point of view of calculation speed, will be most convenient as follows: first one needs to make differentiations with respect to the four momentum variables in (\ref{A2n1}), then make the corresponding changes of the variables following from the delta functions, then sum over the indices $ s_n $ from (\ref{D}), then integrate over the angles in $ d^3 p $, then integrate over imaginary time variables, which requires careful handling of theta functions. The transformations with each of the 24 terms in (\ref{wik}) can be performed independently and using parallel computing tools. We do not give the described intermediate steps, since they are most conveniently performed using the program, and the intermediate formulas themselves are extremely long, while the calculations themselves are not difficult from a mathematical point of view and are done directly. As a result, we obtain the following integral
\begin{eqnarray}
A_2&=&\int_{0}^{\infty} d\tilde{p} e^{\frac{9 \tilde{p}}{2}} \tilde{p}^3 \Big(5600 \tilde{p} \left(49 \tilde{p}^2-95\right) \cosh
   \left(\frac{\tilde{p}}{2}\right)+2016 \tilde{p} \left(25-119 \tilde{p}^2\right) \cosh \left(\frac{3
   \tilde{p}}{2}\right)\nonumber \\
&& +53200 \left(\sinh \left(\frac{3 \tilde{p}}{2}\right)-11 \sinh
   \left(\frac{\tilde{p}}{2}\right)\right) \cosh ^4\left(\frac{\tilde{p}}{2}\right)+\tilde{p} \Big(-224
   \left(\tilde{p}^2+25\right) \cosh \left(\frac{7 \tilde{p}}{2}\right)\nonumber \\
&&+224 \left(119 \tilde{p}^2+575\right) \cosh
   \left(\frac{5 \tilde{p}}{2}\right)+18 \tilde{p} \sinh \left(\frac{\tilde{p}}{2}\right) \Big(-5786
   \tilde{p}^2+\left(\tilde{p}^2+210\right) \cosh \left(3 \tilde{p}\right)\nonumber \\
&&-6 \left(41 \tilde{p}^2+1890\right)
   \cosh \left(2 \tilde{p}\right)+3 \left(1349 \tilde{p}^2+9450\right) \cosh
   \left(\tilde{p}\right)\nonumber \\
&&+39900\Big)\Big)\Big)(50400 \pi ^2 \left(e^{\tilde{p}}+1\right)^9)^{-1}\,,
\label{A2int}
\end{eqnarray}
where the dimensionless variable $ \tilde{p}=|{\bf p}|/T $ was introduced. This integral converges and can be found analytically
\begin{eqnarray}
A_2&=&-\frac{17}{960\pi^2}\,.
\label{A2intres}
\end{eqnarray}
Repeating the entire calculation algorithm for the coefficients $ A_4, A_6 $ in (\ref{AC}), we obtain  at $ m=0 $
\begin{eqnarray}
A_4 = -\frac{17}{2880\pi^2}\,,\quad
A_6 = 0\,.
\label{A4A6}
\end{eqnarray} 
Saving the mass in all formulas, in particular, in the propagators (\ref{prop}), we get more complicated expressions for the coefficients at finite mass given in the Appendix \ref{sec:m}. 

%=================================================
\section{Discussion}
\label{sec:disc}
%=================================================

In the previous section, we described the details of the calculation of the corrections of the fourth order in acceleration to the energy-momentum tensor of the Dirac field, first obtained in \cite{Prokhorov:2019cik}. Taking into account (\ref{A2intres}) and (\ref{A4A6}) we obtain  the next formula for the energy-momentum tensor at $ m=0 $
\begin{eqnarray}
\langle \hat{T}^{\mu\nu}\rangle &=& \Big(\frac{7 \pi ^2 T^4}{60}+\frac{T^2 |a|^2}{24} -\frac{17 |a|^4}{960\pi^2} \Big) u^{\mu}u^{\nu}\nonumber \\
&&- 
\Big(\frac{7 \pi ^2 T^4}{180}+\frac{T^2 |a|^2}{72} -\frac{17|a|^4}{2880\pi^2} \Big) \Delta^{\mu\nu}+\mathcal{O}(a^6)\,,
\label{tensorfermim0}
\end{eqnarray}
where the notation $ |a|=\sqrt{-a_{\mu}a^{\mu}} $ is used.

As discussed in \cite{Becattini:2017ljh, Dowker:1994fi,Frolov:1987dz, Florkowski:2018myy}, the mean value of the energy-momentum tensor calculated in this way should vanish at the proper temperature equal to Unruh temperature. Since the energy-momentum tensor is normalized with respect to the Minkowski vacuum, such a vanishing is a direct consequence of the Unruh effect - an accelerated medium with Unruh temperature corresponds to the Minkowski vacuum. It is easy to see that energy-momentum tensor (\ref{tensorfermim0}) satisfies this condition following from the Unruh effect
\begin{eqnarray}
\langle \hat{T}^{\mu\nu}\rangle  (T=T_U) = 0\,.
\label{UnruhReal}
\end{eqnarray}

Moreover, as discussed in \cite{Prokhorov:2019hif} from the presentation of the result (\ref{tensorfermim0}) in the form of Sommerfeld integrals, as well as comparison with the field theory in a space with a conical singularity, it follows that the calculated fourth order of perturbation theory is maximal, that is, $ \mathcal{O}(a^6)=0 $ at least at $ T>T_U $ \cite{Prokhorov:2019hif}. Thus, Eq.~(\ref{tensorfermim0}) is an exact nonperturbative formula in this region.

We also note that expression (\ref{tensorfermim0}) can be obtained from the point of view of another approach, where field theory in a space with a conical singularity is considered \cite{Dowker:1994fi, Frolov:1987dz}
. As discussed in \cite{Prokhorov:2019hif}, this indicates the duality of the statistical and geometrical approaches to the description of accelerated media.

%=================================================
\section{Conclusions}
\label{sec:concl}
%=================================================

The Zubarev density operator provides a powerful fundamental theoretical method for studying quantum-field effects in the accelerated medium. This makes it possible to obtain information about such a medium from the point of view of an inertial observer and there is no need to go to the curvilinear coordinates of the accelerated frame and consider the features of nontrivial space with a boundary. All effects can be calculated in ordinary flat space described by the Minkowski metric using standard Green functions at finite temperature. In this case, the effects of acceleration are calculated in a regular way in the framework of perturbation theory with the boost operator. However, it is possible to obtain exact nonperturbative expressions in the chiral limit, since the first few orders of the perturbation theory are to give a complete perturbative series.

In particular, earlier in \cite{Prokhorov:2019cik} the Unruh effect for fermions was demonstrated by calculating fourth-order quantum corrections. In the language of the statistical approach with the Zubarev operator, the Unruh effect should lead to the vanishing of the energy-momentum tensor at the proper temperature equal to the Unruh temperature. Thus, the Zubarev density operator allows one to obtain information about the effects associated with the occurrence of an event horizon in an accelerated system and the radiation associated with it.

In this paper, we described the details of the calculations of the coefficients with acceleration in the energy-momentum tensor given in \cite{Prokhorov:2019cik}, focusing on the calculation of the quantum correction to the energy density. The calculation of this correction  consists in finding the mean value of the product of the boost operators and operator of the energy-momentum tensor. Applying Wick theorem, one can transform the average of the product of operators to the product of five thermal propagators. Each of the propagators adds one summation over the Matsubara frequencies and a three-dimensional integral over the momentum, and also each boost operator adds three-dimensional integral over the coordinate and one integral over the imaginary time. The procedure for calculating these sums and integrals is described.  In addition, expressions  for the coefficients at a finite mass are given.

%=================================================
{\bf Acknowledgements}
%=================================================

Useful discussions with F. Becattini are gratefully acknowledged. The work was supported by Russian Science Foundation Grant No 16-12-10059.

\appendix

%=============================================
\section{The coefficients $ a^4 $ at finite mass}
\label{sec:m}
%=============================================

The coefficient $ A_2 $ at a finite mass is described by the expression
\begin{eqnarray}
A_2&=&\int_{0}^{\infty} d\tilde{p}\tilde{p}^2 e^{\frac{9 \tilde{E}_p}{2}} \Big(\tilde{E}_p \Big(9 \left(51450-15619 \tilde{m}^2\right) \tilde{p}^4+175 \left(2450 \tilde{m}^2-361\right) \tilde{p}^2\nonumber \\
&&+175 \tilde{m}^2
   \left(392 \tilde{m}^2-285\right)-140571 \tilde{p}^6\Big) \sinh \left(\frac{\tilde{E}_p}{2}\right)+27 \tilde{E}_p \Big(27 \left(53 \tilde{m}^2+490\right) \tilde{p}^4\nonumber \\
&&+175 \left(70
   \tilde{m}^2-19\right) \tilde{p}^2+35 \tilde{m}^2 \left(56 \tilde{m}^2-75\right)+1431 \tilde{p}^6\Big) \sinh \left(\frac{3 \tilde{E}_p}{2}\right)\nonumber \\
&&-\tilde{E}_p \Big(9 \left(247
   \tilde{m}^2+11550\right) \tilde{p}^4+175 \left(550 \tilde{m}^2+133\right) \tilde{p}^2+175 \tilde{m}^2 \left(88 \tilde{m}^2+105\right)\nonumber \\
&&+2223 \tilde{p}^6\Big) \sinh \left(\frac{5
   \tilde{E}_p}{2}\right)+\tilde{E}_p \Big(9 \left(\tilde{m}^2+210\right) \tilde{p}^4+175 \left(10 \tilde{m}^2+19\right) \tilde{p}^2\nonumber \\
&&+35 \tilde{m}^2 \left(8 \tilde{m}^2+75\right)+9
   \tilde{p}^6\Big) \sinh \left(\frac{7 \tilde{E}_p}{2}\right)-8 \left(\tilde{m}^2+\tilde{p}^2\right) \Big(5 \tilde{m}^2 \left(2 \tilde{p}^2+63\right)\nonumber \\
&&+28 \tilde{p}^2
   \left(\tilde{p}^2+25\right)\Big) \cosh \left(\frac{7 \tilde{E}_p}{2}\right)-504 \left(\tilde{m}^2+\tilde{p}^2\right) \Big(5 \tilde{m}^2 \left(34 \tilde{p}^2-9\right)\nonumber \\
&&+476 \tilde{p}^4-100
   \tilde{p}^2\Big) \cosh \left(\frac{3 \tilde{E}_p}{2}\right)+56 \left(\tilde{m}^2+\tilde{p}^2\right) \Big(5 \tilde{m}^2 \left(34 \tilde{p}^2+207\right)\nonumber \\
&&+476 \tilde{p}^4+2300
   \tilde{p}^2\Big) \cosh \left(\frac{5 \tilde{E}_p}{2}\right)+1400 \left(\tilde{m}^2+\tilde{p}^2\right) \Big(\tilde{m}^2 \left(70 \tilde{p}^2-171\right)\nonumber \\
&&+4 \tilde{p}^2 \left(49
   \tilde{p}^2-95\right)\Big) \cosh \left(\frac{\tilde{E}_p}{2}\right)\Big)(50400 \pi ^2 \left(e^{\tilde{E}_p}+1\right){}^9 \tilde{E}_p^2)^{-1}\,,
\label{A2intm}
\end{eqnarray}
where the dimensionless quantities $ \tilde{m}=m/T, \tilde{E}_p=\sqrt{p^2+m^2}/T $ are introduced. The coefficient $ A_4 $ at a finite mass has the form
\begin{eqnarray}
A_4&=&\int_{0}^{\infty} d\tilde{p}\Big(\tilde{p}^4 e^{\frac{9 \tilde{\text{E}}_p}{2}} \Big(1960 \sinh \left(\frac{\tilde{\text{E}}_p}{2}\right) \cosh ^2\left(\frac{\tilde{\text{E}}_p}{2}\right) \Big(\left(8 \tilde{m}^2+15\right)
   \cosh \left(2 \tilde{\text{E}}_p\right)\nonumber \\
&&-8 \left(56 \tilde{m}^2+15\right) \cosh \left(\tilde{\text{E}}_p\right)+984 \tilde{m}^2-135\Big)+\tilde{p}^2 \left(408170-421713 \tilde{p}^2\right) \sinh
   \left(\frac{\tilde{\text{E}}_p}{2}\right)\nonumber \\
&&+27 \tilde{p}^2 \left(4293 \tilde{p}^2+11662\right) \sinh \left(\frac{3 \tilde{\text{E}}_p}{2}\right)-\tilde{p}^2 \left(6669 \tilde{p}^2+91630\right)
   \sinh \left(\frac{5 \tilde{\text{E}}_p}{2}\right)\nonumber \\
&&+\tilde{p}^2 \left(27 \tilde{p}^2+1666\right) \sinh \left(\frac{7 \tilde{\text{E}}_p}{2}\right)+29400 \left(14 \tilde{p}^2-19\right)
   \tilde{\text{E}}_p \cosh \left(\frac{\tilde{\text{E}}_p}{2}\right)\nonumber \\
&&-168 \left(2 \tilde{p}^2+35\right) \tilde{\text{E}}_p \cosh \left(\frac{7 \tilde{\text{E}}_p}{2}\right)-10584 \left(34
   \tilde{p}^2-5\right) \tilde{\text{E}}_p \cosh \left(\frac{3 \tilde{\text{E}}_p}{2}\right)\nonumber \\
&&+1176 \left(34 \tilde{p}^2+115\right) \tilde{\text{E}}_p \cosh \left(\frac{5
   \tilde{\text{E}}_p}{2}\right)\Big)\Big)\left(1058400 \pi ^2 \left(e^{\tilde{\text{E}}_p}+1\right){}^9 \tilde{\text{E}}_p\right)^{-1}\,.
\label{A4intm}
\end{eqnarray}
The coefficient $ A_6 $ is zero both for massless and massive Dirac fields
\begin{eqnarray}
A_6 &=& 0\,.
\label{A6intm}
\end{eqnarray}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%% optional
%\sampleavailability{Samples of the compounds ...... are available from the authors.}

%% for journal Sci
%\reviewreports{\\
%Reviewer 1 comments and authors’ response\\
%Reviewer 2 comments and authors’ response\\
%Reviewer 3 comments and authors’ response
%}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\end{document}